OpenStudy (destinyyyy):

Can someone check my work.. Find the coordinates of vertex, y-intercept, x-intercept (if any exist).

OpenStudy (destinyyyy):

g(x)= -x^2 -6x+40

OpenStudy (destinyyyy):

-(-6)/2(-1) = 6/-2 =3 g(3)= -(3)^2 -6(3)+40 31 vertex- (3,31)

OpenStudy (destinyyyy):

I see a mistake.. -3

OpenStudy (sdfgsdfgs):

-3 is the right x for vertex :) wats the y then?

OpenStudy (destinyyyy):

Vertex- (-3,49)

OpenStudy (destinyyyy):

y-inter- (0,40)

OpenStudy (sdfgsdfgs):

both correct!

OpenStudy (destinyyyy):

x-inter- (-10,0),(4,0)

OpenStudy (sdfgsdfgs):

correct as well! u can click on the "properties" button the next time: https://www.wolframalpha.com/input/?i=y%3D-x%5E2+-6x%2B40

OpenStudy (destinyyyy):

Do you mind checking another one?

OpenStudy (sdfgsdfgs):

@Destinyyyy hahahaaa do u wanna give wolframalpha a try first? i can take a look if u dont like the wolf :)

OpenStudy (destinyyyy):

I dont really know how to use that and I dont entirely trust those kind of sites. I use mathway and its checked my answers wrong before..

OpenStudy (destinyyyy):

If you dont want to check it, its fine

OpenStudy (sdfgsdfgs):

OK sure i will do it :)

OpenStudy (destinyyyy):

Okay thank you.. f(x)=x^2+6x+8 vertex- (-3,-1) y-inter- (0,8) x-inter-(-2,0),(-4,0)

OpenStudy (sdfgsdfgs):

@Destinyyyy all correct! GREAT job :)

OpenStudy (destinyyyy):

Thanks..

OpenStudy (sdfgsdfgs):

welcome